# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

### v3

@@ 8 Paraphrased sentence:
(±0.005)
0.309
(±0.007)
0.753 0.334
gemma-7b 0.501 0.404
0.484
(±0.006)
0.568
(±0.009)
0.581
(±0.008)
0.638
(±0.008)
0.630
(±0.002)
0.675
(±0.006)
0.803 0.709
gemma-1.1-
7b-it
0.656 0.614
0.680
(±0.006)
0.627
(±0.009)
0.728
(±0.006)
0.657
(±0.006)
0.748
(±0.005)
0.668
(±0.006)
0.775 0.692
Llama-3.2-3B 0.416 0.324
0.471
(±0.006)
0.381
(±0.008)
0.561
(±0.007)
0.451
(±0.008)
0.590
(±0.008)
0.491
(±0.009)
0.754 0.282
Llama-3.2-3B-
Instruct
0.454 0.516
0.477
(±0.006)
0.464
(±0.010)
0.546
(±0.005)
0.452
(±0.008)
0.589
(±0.004)
0.450
(±0.006)
0.769 0.386
Llama-3.1-8B 0.462 0.477
0.499
(±0.005)
0.512
(±0.013)
0.598
(±0.006)
0.591
(±0.008)
0.645
(±0.005)
0.630
(±0.009)
0.817 0.608
Llama-3.1-8B-
Instruct
0.519 0.684
0.557
(±0.005)
0.640
(±0.004)
0.628
(±0.002)
0.663
(±0.004)
0.665
(±0.007)
0.681
(±0.014)
0.784 0.762
Mistral-7B-v0.3 0.446 0.297
0.475
(±0.006)
0.571
(±0.013)
0.560
(±0.011)
0.683
(±0.010)
0.608
(±0.002)
0.694
(±0.007)
0.781 0.427
Mistral-7B-
Instruct-v0.3
0.596 0.630
0.635
(±0.006)
0.622
(±0.007)
0.685
(±0.005)
0.630
(±0.011)
0.707
(±0.009)
0.632
(±0.007)
0.765 0.629
Qwen2.5-3B 0.427 0.531
0.474
(±0.007)
0.490
(±0.007)
0.540
(±0.003)
0.499
(±0.006)
0.574
(±0.007)
0.500
(±0.007)
0.722 0.648
Qwen2.5-3B-
Instruct
0.474 0.585
0.530
(±0.008)
0.511
(±0.006)
0.628
(±0.007)
0.512
(±0.006)
0.665
(±0.010)
0.513
(±0.006)
0.696 0.713
Qwen2.5-7B 0.454 0.628
0.483
(±0.009)
0.606
(±0.005)
0.541
(±0.007)
0.622
(±0.007)
0.587
(±0.002)
0.639
(±0.007)
0.745 0.706
Qwen2.5-7B-
Instruct
0.536 0.636
0.549
(±0.007)
0.634
(±0.009)
0.634
(±0.010)
0.634
(±0.008)
0.674
(±0.006)
0.636
(±0.004)
0.734 0.764
gpt-4o-mini - 0.621 -
0.659
(±0.008)
0.688
(±0.005)
0.700
(±0.009)
gpt-4.1-mini - 0.842 -
0.845
(±0.007)
0.857
(±0.005)
0.867
(±0.005)
claude-3-5-
haiku-latest
- 0.688 -
0.495
(±0.006)
0.470
(±0.008)
0.613
(±0.013)
2-3B models
average
0.436 0.399 0.469 0.405 0.555 0.418 0.591 0.422 0.736 0.437
7-8B models
average
0.521 0.546 0.545 0.598 0.619 0.640 0.658 0.657 0.775 0.662
api models
average
- 0.717 - 0.666 - 0.672 - 0.727 - -
base models
average
0.450 0.405 0.473 0.484 0.557 0.537 0.599 0.556 0.763 0.520
instruct-tuned
models average
0.520 0.561 0.552 0.546 0.627 0.552 0.660 0.556 0.754 0.611
total average 0.485 0.524 0.513 0.542 0.592 0.567 0.629 0.586 0.758 0.566
Table 12: Zero-shot, few-shot, and SFT evaluation results on Thunder-NUBench with the definition prompt. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column (excluding API models).
zero-shot 1-shot 5-shot 10-shot After SFT
evaluation
setting
comple-
tion
option
comple-
tion
(±SD)
option
(±SD)
comple-
tion
(±SD)
option
(±SD)
comple-
tion
(±SD)
option
(±SD)
comple-
tion
option
gemma-2b 0.404 0.257
0.423
(±0.003)
0.253
(±0.005)
0.511
(±0.003)
0.257
(±0.012)
0.547
(±0.004)
0.145
(±0.009)
0.496 0.514
gemma-1.1-
2b-it
0.394 0.260
0.440
(±0.007)
0.282
(±0.005)
0.524
(±0.009)
0.241
(±0.007)
0.561
(±0.009)
0.242
(±0.009)
0.519 0.476
gemma-7b 0.462 0.410
0.472
(±0.005)
0.483
(±0.008)
0.564
(±0.004)
0.531
(±0.007)
0.607
(±0.010)
0.549
(±0.008)
0.731 0.635
gemma-1.1-
7b-it
0.588 0.472
0.653
(±0.008)
0.510
(±0.009)
0.708
(±0.009)
0.501
(±0.004)
0.728
(±0.006)
0.533
(±0.011)
0.725 0.616
Llama-3.2-3B 0.445 0.306
0.463
(±0.005)
0.317
(±0.011)
0.550
(±0.008)
0.365
(±0.010)
0.582
(±0.006)
0.407
(±0.008)
0.567 0.553
Llama-3.2-3B-
Instruct
0.454 0.429
0.440
(±0.006)
0.387
(±0.009)
0.501
(±0.007)
0.381
(±0.008)
0.560
(±0.005)
0.389
(±0.006)
0.586 0.595
Llama-3.1-8B 0.442 0.402
0.466
(±0.004)
0.394
(±0.008)
0.566
(±0.006)
0.462
(±0.012)
0.624
(±0.005)
0.518
(±0.010)
0.726 0.642
Llama-3.1-8B-
Instruct
0.508 0.539
0.529
(±0.008)
0.515
(±0.010)
0.627
(±0.005)
0.554
(±0.004)
0.669
(±0.003)
0.580
(±0.006)
0.673 0.681
Mistral-7B-v0.3 0.454 0.236
0.485
(±0.004)
0.416
(±0.014)
0.551
(±0.010)
0.597
(±0.009)
0.591
(±0.007)
0.631
(±0.007)
0.586 0.545
Mistral-7B-
Instruct-v0.3
0.588 0.586
0.608
(±0.013)
0.551
(±0.010)
0.662
(±0.005)
0.556
(±0.008)
0.685
(±0.005)
0.558
(±0.005)
0.672 0.646
Qwen2.5-3B 0.415 0.463
0.461
(±0.005)
0.399
(±0.002)
0.532
(±0.005)
0.324
@@
0.662 0.615
Qwen2.5-3B-
@@
0.467 0.548
@@
0.691 0.689
Qwen2.5-7B-
@@
0.507 0.545
0.511
(±0.006)
0.527
(±0.008)
0.621
(±0.012)
0.526
(±0.007)
0.662
(±0.004)
0.544
(±0.003)
0.674 0.705
gpt-4o-mini - 0.595 -
0.620
(±0.014)
0.647
(±0.003)
0.669
(±0.006)
gpt-4.1-mini - 0.858 -
0.864
(±0.004)
0.876
(±0.003)
0.870
(±0.005)
claude-3-5-
haiku-latest
- 0.604 -
0.472
(±0.009)
0.537
(±0.009)
0.689
(±0.013)
2-3B models
average
0.430 0.377 0.453 0.352 0.545 0.299 0.582 0.332 0.584 0.554
7-8B models
average
0.499 0.468 0.524 0.489 0.605 0.534 0.643 0.560 0.685 0.645
api models
average
- 0.686 - 0.652 - 0.687 - 0.743 - -
base models
average
0.438 0.376 0.462 0.397 0.544 0.440 0.585 0.447 0.637 0.599
instruct-tuned
models average
0.501 0.483 0.525 0.464 0.614 0.427 0.648 0.477 0.646 0.613
total average 0.469 0.474 0.493 0.469 0.579 0.478 0.617 0.512 0.642 0.606
Table 13: Zero-shot, few-shot, and SFT evaluation results on Thunder-NUBench with the detail prompt. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column (excluding API models).
Following the classification of prompts introduced in Appendix O, we report results on Thunder-NUBench separately for the simple, definition, and detail prompt styles. The experiments are conducted under three evaluation regimes: (i) zero-shot, (ii) few-shot (1-, 5-, and 10-shot, averaged across seeds), and (iii) supervised fine-tuning (SFT). For SFT, we assess models only in the zero-shot setting in order to directly measure the effect of task-specific training without the influence of in-context examples.
Figure 4: Model performance on Thunder-NUBench with definition prompt. Circles (blue) represent the average performance of 2-3B models, squares (purple) indicate the average for 7-8B models, upward triangles (orange) signify the average of base models, and downward triangles (red) denote the average of instruction-tuned models. Stars (green) represent API models.
Figure 5: Model performance on Thunder-NUBench with detail prompt. Circles (blue) represent the average performance of 2-3B models, squares (purple) indicate the average for 7-8B models, upward triangles (orange) signify the average of base models, and downward triangles (red) denote the average of instruction-tuned models. Stars (green) represent API models.
Table 11, Table 12, and Table 13 present the complete numerical results for the simple, definition, and detail prompts, respectively. The same results are also summarized in graphical form in Figure 3, Figure 4, and Figure 5.
Across prompt types, we observe the following trends:
  • Simple prompts. Results for simple prompts are described in Section 5.2; we summarize them here only for completeness.
  • Definition prompts. Performance patterns closely resemble those of simple prompts. However, in the option-selection evaluation, 2-3B models show smaller gains from additional shots, while 7-8B models tend to achieve slightly higher accuracy after SFT compared to 10-shot prompting.
  • Detail prompts. For smaller (2-3B) models in the option-selection evaluation, adding more shots can actually reduce performance, reflecting difficulties with long and complex prompt instructions to these small models. Nonetheless, these models, along with larger ones, show larger SFT gains under detail prompts than under simple or definition prompts.
Overall, these findings show that prompt formulation plays a non-trivial role in shaping performance trends. In all cases, SFT consistently boosts zero-shot performance.
Appendix Q General Benchmark Performance after SFT
tasks
arc_
challenge
arc_
easy
gsm8k hellaswag
openbook-
qa
wino-
grande
avg
metric acc acc
exact_
match
acc_norm acc_norm acc
Before SFT with simple prompt on Thunder-NUBench
2-3B models
average
0.432 0.747 0.325 0.714 0.409 0.668 0.549
7-8B models
average
0.514 0.810 0.592 0.797 0.458 0.727 0.65
base models
average
0.464 0.784 0.480 0.768 0.437 0.709 0.607
instruct-tuned
models average
0.493 0.782 0.475 0.756 0.437 0.694 0.606
total average 0.479 0.783 0.478 0.762 0.437 0.702 0.607
After SFT with simple prompt on Thunder-NUBench
2-3B models
average
0.431 0.746 0.365 0.713 0.411 0.671 0.556
7-8B models
average
0.524 0.819 0.536 0.798 0.453 0.721 0.642
base models
average
0.474 0.788 0.428 0.772 0.437 0.706 0.601
instruct-tuned
models average
0.495 0.787 0.497 0.751 0.432 0.693 0.609
total average 0.484 0.788 0.463 0.762 0.435 0.699 0.605
Before SFT with definition prompt on Thunder-NUBench
2-3B models
average
0.431 0.747 0.325 0.714 0.409 0.668 0.549
7-8B models
average
0.514 0.810 0.592 0.797 0.458 0.727 0.65
base models
average
0.464 0.784 0.480 0.768 0.437 0.709 0.607
instruct-tuned
models average
0.493 0.782 0.475 0.755 0.437 0.694 0.606
total average 0.479 0.783 0.478 0.762 0.437 0.701 0.607
After SFT with definition prompt on Thunder-NUBench
2-3B models
average
0.433 0.745 0.341 0.713 0.407 0.669 0.551
7-8B models
average
0.528 0.819 0.545 0.799 0.455 0.725 0.645
base models
average
0.477 0.785 0.419 0.772 0.436 0.709 0.600
instruct-tuned
models average
0.498 0.790 0.496 0.753 0.433 0.694 0.611
total average 0.487 0.788 0.457 0.762 0.434 0.701 0.605
Before SFT with detail prompt on Thunder-NUBench
2-3B models
average
0.431 0.747 0.325 0.714 0.409 0.668 0.549
7-8B models
average
0.514 0.810 0.592 0.797 0.458 0.727 0.65
base models
average
0.464 0.784 0.480 0.768 0.437 0.709 0.607
instruct-tuned
models average
0.493 0.782 0.475 0.755 0.437 0.694 0.606
total average 0.479 0.783 0.478 0.762 0.437 0.701 0.607
After SFT with detail prompt on Thunder-NUBench
2-3B models
average
0.429 0.746 0.359 0.713 0.409 0.668 0.554
7-8B models
average
0.523 0.816 0.551 0.8 0.458 0.723 0.645
base models
average
0.469 0.785 0.424 0.773 0.439 0.704 0.599
instruct-tuned
models average
0.496 0.787 0.513 0.753 0.434 0.695 0.613
total average 0.483 0.786 0.469 0.763 0.437 0.699 0.606
Table 14: Average performance of 2–3B / 7–8B, and base / instruction-tuned models on six general benchmarks before and after SFT on Thunder-NUBench. Here, acc denotes accuracy, acc_norm is normalized accuracy, and exact_match requires an exact string match with the reference answer.
To assess whether supervised fine-tuning (SFT) on Thunder-NUBench affects performance on broader natural language understanding tasks, we evaluate models on six widely used benchmarks: ARC-Challenge, ARC-Easy, GSM8K, HellaSwag, OpenBookQA, and WinoGrande (clark2018think; cobbe2021training; zellers2019hellaswag; mihaylov2018can; sakaguchi2021winogrande). These datasets cover a diverse range of domains, including commonsense reasoning, scientific knowledge, mathematics, and reading comprehension.
Following the classification of prompts introduced in Appendix O, SFT is conducted separately for the simple, definition, and detail prompt styles. We then report post-SFT zero-shot performance across the six benchmarks to verify whether task-specific training on Thunder-NUBench preserves general capabilities.
Table 14 summarizes the results. We find that performance on general benchmarks remains broadly stable after SFT, indicating that fine-tuning on Thunder-NUBench does not substantially harm general capabilities.
Appendix R Total Analysis of Model Predictions on the Thunder-NUBench
This appendix extends the error analysis presented in Section 5.3, providing the complete results. In particular, we examine (i) incorrect choice distributions and (ii) confusion rates for local negation categories, comparing models of different sizes (2-3B vs. 7-8B) and training paradigms (pretrained vs. instruction-tuned) under both zero-shot and few-shot conditions.
We report few-shot results using a fixed random seed (1234), which corresponds to the default seed used in the LM Evaluation Harness framework. Averaging over multiple seeds was avoided, as it could obscure specific error patterns and make fine-grained confusion analysis less interpretable.
For consistency with the main text, we report results only under the simple prompt, which serves as the default evaluation setting throughout the paper.
We organize the complete prediction analysis by model family.
  • Results for the Gemma family are reported in Table 15, Table 16.
  • Results for the LLaMA family are reported in Table 17, Table 18.
  • Results for the Mistral family are reported in Table 19, Table 20.
  • Results for the Qwen family are reported in Table 21, Table 22.
  • Results for API models are reported in Table 23.
Each table follows the same format, reporting error rates, incorrect choice distributions, and confusion rates across local negation subcategories under zero-shot, few-shot, and SFT conditions.
In the option-selection evaluation setting, we also track cases labeled as “Answer Format Wrong.” This category captures instances where the model’s output does not follow the required answer format (selecting strictly one of A, B, C, or D). Because such responses cannot be mapped to a specific incorrect option, they are not included in the incorrect choice distribution or confusion rate. Instead, we report their raw counts alongside the other error statistics. This also serves as an indicator of the model’s ability to follow output-format instructions.
Overall, we observe consistent trends that larger models (7-8B) achieve lower error rates than smaller ones (2-3B), and introduction-tuned variants generally outperform base models. Supervised fine-tuning (SFT) tends to reduce error rates. Meanwhile, incorrect choice distributions and local negation confusion rates vary across models and settings, showing that this analysis can serve as a useful tool to identify which aspects of negation remain particularly challenging.
Notably, we also find some unusual behaviors. For example, the Gemma-1.1-2B-it model after SFT produces 1,239 outputs with incorrect answer formats in the option-selection setting, indicating formatting issues rather than genuine reasoning errors. Similarly, for API models, Claude 3.5 Haiku shows increasing formatting errors as more shots are added, which directly degrades performance.
Error
Rate
Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)

### v4

@@ 8 Paraphrased sentence:
0.715
(±0.009)
0.764 0.805
Qwen3-0.6B-Base 0.424 0.342
0.487
(±0.008)
0.397
(±0.004)
0.564
(±0.006)
0.416
(±0.008)
0.607
(±0.006)
0.461
(±0.010)
0.723 0.325
Qwen3-0.6B 0.413 0.459
0.411
(±0.004)
0.414
(±0.009)
0.487
(±0.003)
0.468
(±0.003)
0.535
(±0.007)
0.495
(±0.008)
0.679 0.526
Qwen3-1.7B-Base 0.462 0.483
0.511
(±0.007)
0.542
(±0.004)
0.574
(±0.005)
0.620
(±0.002)
0.597
(±0.005)
0.653
(±0.003)
0.724 0.609
Qwen3-1.7B 0.420 0.447
0.442
(±0.008)
0.502
(±0.013)
0.490
(±0.006)
0.547
(±0.007)
0.522
(±0.006)
0.556
(±0.010)
0.642 0.573
Qwen3-4B-Base 0.485 0.585
0.503
(±0.005)
0.581
(±0.006)
0.546
(±0.004)
0.636
(±0.006)
0.566
(±0.006)
0.649
(±0.009)
0.729 0.711
Qwen3-4B 0.439 0.577
0.489
(±0.010)
0.635
(±0.003)
0.559
(±0.008)
0.652
(±0.004)
0.600
(±0.005)
0.685
(±0.003)
0.688 0.711
Qwen3-8B-Base 0.455 0.668
0.513
(±0.004)
0.700
(±0.007)
0.557
(±0.005)
0.737
(±0.004)
0.591
(±0.009)
0.755
(±0.009)
0.714 0.823
Qwen3-8B 0.473 0.619
0.498
(±0.014)
0.695
(±0.006)
0.548
(±0.007)
0.693
(±0.007)
0.586
(±0.007)
0.723
(±0.007)
0.740 0.760
Qwen3-14B-Base 0.508 0.736
0.553
(±0.008)
0.771
(±0.010)
0.628
(±0.004)
0.797
(±0.007)
0.660
(±0.007)
0.816
(±0.006)
Qwen3-14B 0.488 0.720
0.522
(±0.007)
0.718
(±0.011)
0.601
(±0.008)
0.762
(±0.009)
0.636
(±0.005)
0.766
(±0.005)
Qwen3-32B 0.514 0.761
0.570
(±0.011)
0.814
(±0.004)
0.657
(±0.003)
0.808
(±0.001)
0.695
(±0.004)
0.821
(±0.005)
Llama-3.1-8B 0.451 0.431
0.486
(±0.006)
0.484
(±0.003)
0.564
(±0.008)
0.560
(±0.010)
0.620
(±0.004)
0.611
(±0.010)
0.802 0.672
Llama-3.1-8B-
@@
0.532 0.514
0.589
(±0.004)
0.588
(±0.010)
0.643
(±0.007)
0.626
(±0.013)
0.679
(±0.007)
0.648
(±0.004)
0.789 0.730
Llama-3.2-1B 0.403 0.259
0.439
(±0.009)
0.263
(±0.007)
0.497
(±0.007)
0.249
(±0.009)
0.527
(±0.005)
0.264
(±0.008)
0.747 0.259
Llama-3.2-1B-
Instruct
0.416 0.300
0.468
(±0.003)
0.347
(±0.009)
0.523
(±0.008)
0.339
(±0.010)
0.542
(±0.009)
0.330
(±0.009)
0.703 0.263
Llama-3.2-3B 0.447 0.374
0.478
(±0.003)
0.419
(±0.006)
0.568
(±0.006)
0.470
(±0.006)
0.590
(±0.004)
0.479
(±0.008)
0.761 0.364
Llama-3.2-3B-
Instruct
0.488 0.488
0.464
(±0.010)
0.518
(±0.014)
0.540
(±0.007)
@@
0.656
(±0.005)
0.786
(±0.005)
Mistral-Small-24B-
Instruct-2501
0.581 0.729
0.638
(±0.012)
0.758
(±0.007)
0.668
(±0.006)
0.757
(±0.007)
0.702
(±0.005)
0.804
(±0.005)
total average 0.481 0.520 0.513 0.557 0.580 0.591 0.615 0.612 0.740 0.564
Table 15: Zero-shot, few-shot, and SFT evaluation results on Thunder-NUBench with the detailed instruction. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column.
models zero-shot 1-shot (±SD) 5-shot (±SD) 10-shot (±SD)
GPT-4o mini 0.713 0.737 (±0.011) 0.773 (±0.005) 0.795 (±0.011)
GPT-4o 0.782 0.785 (±0.005) 0.800 (±0.012) 0.824 (±0.006)
GPT-4.1 mini 0.757 0.800 (±0.006) 0.840 (±0.007) 0.851 (±0.007)
GPT-4.1 0.881 0.863 (±0.003) 0.874 (±0.003) 0.891 (±0.001)
Haiku 4.5 0.772 0.775 (±0.005) 0.816 (±0.008) 0.837 (±0.006)
Sonnet 4.5 0.876 0.880 (±0.006) 0.875 (±0.002) 0.893 (±0.009)
total average 0.797 0.807 0.830 0.849
Table 16: Zero-shot and few-shot results of API models on Thunder-NUBench with the definition instruction. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column.
models zero-shot 1-shot(±SD) 5-shot (±SD) 10-shot (±SD)
GPT-4o mini 0.756 0.768 (±0.007) 0.774 (±0.005) 0.790 (±0.011)
GPT-4o 0.863 0.847 (±0.006) 0.825 (±0.012) 0.826 (±0.006)
GPT-4.1 mini 0.902 0.873 (±0.004) 0.870 (±0.007) 0.870 (±0.007)
GPT-4.1 0.936 0.915 (±0.004) 0.901 (±0.003) 0.899 (±0.001)
Haiku 4.5 0.865 0.870 (±0.004) 0.870 (±0.008) 0.879 (±0.006)
Sonnet 4.5 0.915 0.914 (±0.004) 0.882 (±0.002) 0.825 (±0.009)
total average 0.873 0.865 0.854 0.848
Table 17: Zero-shot and few-shot results of API models on Thunder-NUBench with the detailed instruction. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column.
tasks ARC-Challenge ARC-Easy GSM8K HellaSwag OpenBookQA WinoGrande average
metric acc acc exact_match acc_norm acc_norm acc
gemma-2-2b 0.468 (+0.001) 0.797 (- 0.002) 0.259 (+0.002) 0.740 (- 0.001) 0.422 (- 0.002) 0.680 (+0.002) 0.561 (+0.000)
gemma-2-2b-it 0.498 (+0.003) 0.809 (+0.002) 0.438 (- 0.014) 0.716 (- 0.001) 0.434 (- 0.002) 0.680 (+0.007) 0.596 (- 0.001)
gemma-2-9b 0.616 (+0.005) 0.874 (+0.002) 0.672 (+0.035) 0.800 (+0.012) 0.472 (- 0.006) 0.741 (+0.009) 0.696 (+0.010)
gemma-2-9b-it 0.629 (+0.008) 0.858 (+0.010) 0.820 (+0.013) 0.800 (+0.000) 0.502 (- 0.008) 0.762 (- 0.010) 0.729 (+0.002)
Qwen3-0.6B-Base 0.326 (- 0.006) 0.668 (- 0.007) 0.510 (- 0.005) 0.517 (+0.002) 0.340 (+0.000) 0.594 (- 0.009) 0.493 (- 0.004)
Qwen3-0.6B 0.317 (- 0.010) 0.607 (- 0.011) 0.409 (+0.019) 0.473 (- 0.009) 0.318 (- 0.006) 0.560 (- 0.006) 0.447 (- 0.004)
Qwen3-1.7B-Base 0.416 (- 0.003) 0.733 (+0.013) 0.723 (- 0.034) 0.665 (+0.000) 0.392 (- 0.002) 0.639 (+0.006) 0.595 (- 0.003)
Qwen3-1.7B 0.403 (+0.009) 0.726 (+0.017) 0.692 (- 0.012) 0.604 (+0.004) 0.366 (- 0.004) 0.608 (+0.010) 0.567 (+0.004)
Qwen3-4B-Base 0.477 (+0.016) 0.789 (+0.006) 0.842 (- 0.008) 0.737 (+0.004) 0.408 (- 0.010) 0.706 (- 0.002) 0.660 (+0.001)
Qwen3-4B 0.507 (+0.017) 0.804 (+0.012) 0.840 (+0.026) 0.684 (+0.003) 0.402 (- 0.004) 0.661 (+0.004) 0.650 (+0.010)
Qwen3-8B-Base 0.532 (+0.002) 0.817 (+0.005) 0.850 (- 0.017) 0.785 (+0.003) 0.410 (+0.012) 0.723 (+0.002) 0.686 (+0.001)
Qwen3-8B 0.555 (+0.008) 0.835 (+0.005) 0.880 (+0.002) 0.750 (+0.006) 0.416 (+0.010) 0.677 (+0.017) 0.686 (+0.008)
Llama-3.1-8B 0.513 (+0.015) 0.814 (+0.014) 0.502 (+0.027) 0.790 (+0.007) 0.450 (+0.016) 0.739 (+0.004) 0.635 (+0.014)
Llama-3.1-8B-
@@
0.515 (+0.015) 0.819 (+0.006) 0.784 (- 0.012) 0.792 (- 0.004) 0.432 (+0.008) 0.735 (+0.004) 0.680 (+0.003)
Llama-3.2-1B 0.317 (- 0.012) 0.654 (- 0.010) 0.068 (- 0.011) 0.636 (- 0.004) 0.372 (+0.000) 0.603 (+0.010) 0.442 (- 0.005)
Llama-3.2-1B-
Instruct
0.358 (- 0.007) 0.683 (+0.002) 0.334 (- 0.002) 0.607 (- 0.006) 0.342 (+0.006) 0.597 (+0.004) 0.487 (- 0.001)
Llama-3.2-3B 0.422 (- 0.007) 0.743 (- 0.003) 0.262 (+0.025) 0.736 (+0.007) 0.428 (- 0.004) 0.698 (- 0.005) 0.548 (+0.002)
Llama-3.2-3B-
Instruct
0.436 (+0.003) 0.741 (+0.007) 0.649 (- 0.008) 0.705 (- 0.003) 0.362 (+0.002) 0.678 (- 0.001) 0.595 (+0.000)
Mistral-7B-v0.3 0.488 (+0.015) 0.796 (+0.011) 0.371 (- 0.053) 0.804 (+0.009) 0.436 (- 0.006) 0.736 (- 0.002) 0.605 (- 0.004)
Mistral-7B-Instruct-
v0.3
0.572 (- 0.008) 0.842 (- 0.007) 0.498 (- 0.043) 0.829 (- 0.004) 0.474 (- 0.016) 0.744 (- 0.014) 0.660 (- 0.015)
total average 0.601 (+0.001)
Table 18: General benchmark performance after supervised fine-tuning (SFT) on Thunder-NUBench with the definition instruction. Here, acc denotes accuracy, acc_norm is normalized accuracy, and exact_match requires an exact string match with the reference answer. For each model and task, the main value reports the performance of the pre-SFT model, while the value in parentheses indicates the change after applying SFT.
tasks ARC-Challenge ARC-Easy GSM8K HellaSwag OpenBookQA WinoGrande average
metric acc acc exact_match acc_norm acc_norm acc
gemma-2-2b 0.468 (+0.000) 0.797 (+0.000) 0.259 (+0.000) 0.740 (+0.000) 0.422 (+0.000) 0.680 (+0.000) 0.561 (+0.000)
gemma-2-2b-it 0.498 (- 0.001) 0.809 (- 0.002) 0.438 (+0.000) 0.716 (- 0.001) 0.434 (+0.002) 0.680 (+0.002) 0.596 (+0.000)
gemma-2-9b 0.616 (+0.004) 0.874 (+0.001) 0.672 (+0.035) 0.800 (+0.013) 0.472 (- 0.008) 0.741 (+0.006) 0.696 (+0.009)
gemma-2-9b-it 0.629 (+0.006) 0.858 (+0.011) 0.820 (+0.005) 0.800 (+0.001) 0.502 (+0.002) 0.762 (- 0.004) 0.729 (+0.004)
Qwen3-0.6B-Base 0.326 (+0.000) 0.668 (+0.000) 0.510 (+0.000) 0.517 (+0.000) 0.340 (+0.000) 0.594 (+0.000) 0.493 (+0.000)
Qwen3-0.6B 0.317 (- 0.009) 0.607 (+0.016) 0.409 (+0.014) 0.473 (- 0.012) 0.318 (- 0.002) 0.560 (- 0.001) 0.447 (+0.001)
Qwen3-1.7B-Base 0.416 (+0.002) 0.733 (+0.003) 0.723 (- 0.035) 0.665 (+0.000) 0.392 (- 0.004) 0.639 (+0.002) 0.595 (- 0.005)
Qwen3-1.7B 0.403 (+0.003) 0.726 (+0.021) 0.692 (- 0.024) 0.604 (+0.003) 0.366 (+0.000) 0.608 (+0.009) 0.567 (+0.002)
Qwen3-4B-Base 0.477 (+0.026) 0.789 (+0.009) 0.842 (- 0.036) 0.737 (+0.005) 0.408 (- 0.004) 0.706 (+0.004) 0.660 (+0.001)
Qwen3-4B 0.507 (+0.008) 0.804 (+0.009) 0.840 (+0.008) 0.684 (+0.008) 0.402 (+0.006) 0.661 (+0.006) 0.650 (+0.008)
Qwen3-8B-Base 0.532 (+0.024) 0.817 (+0.011) 0.850 (- 0.040) 0.785 (+0.004) 0.410 (+0.014) 0.723 (- 0.002) 0.686 (+0.002)
Qwen3-8B 0.555 (+0.002) 0.835 (+0.000) 0.880 (- 0.002) 0.750 (+0.005) 0.416 (+0.006) 0.677 (+0.016) 0.686 (+0.005)
Llama-3.1-8B 0.513 (+0.003) 0.814 (+0.012) 0.502 (- 0.001) 0.790 (+0.010) 0.450 (+0.018) 0.739 (+0.001) 0.635 (+0.007)
Llama-3.1-8B-
Instruct
0.515 (+0.009) 0.819 (+0.003) 0.784 (- 0.013) 0.792 (- 0.006) 0.432 (+0.006) 0.735 (+0.009) 0.680 (+0.001)
Llama-3.2-1B 0.317 (- 0.009) 0.654 (- 0.005) 0.068 (- 0.024) 0.636 (- 0.008) 0.372 (- 0.004) 0.603 (+0.017) 0.442 (- 0.006)
Llama-3.2-1B-
Instruct
0.358 (- 0.004) 0.683 (+0.002) 0.334 (- 0.008) 0.607 (- 0.007) 0.342 (+0.008) 0.597 (+0.000) 0.487 (- 0.002)
Llama-3.2-3B 0.422 (- 0.005) 0.743 (- 0.003) 0.262 (+0.014) 0.736 (+0.004) 0.428 (- 0.006) 0.698 (+0.004) 0.548 (+0.001)
Llama-3.2-3B-
Instruct
0.436 (+0.001) 0.741 (+0.006) 0.649 (+0.001) 0.705 (- 0.001) 0.362 (+0.000) 0.678 (+0.000) 0.595 (+0.001)
Mistral-7B-v0.3 0.488 (+0.013) 0.796 (+0.011) 0.371 (- 0.094) 0.804 (+0.004) 0.436 (+0.008) 0.736 (+0.005) 0.605 (- 0.009)
Mistral-7B-Instruct-
v0.3
0.572 (- 0.009) 0.842 (- 0.004) 0.498 (- 0.057) 0.829 (- 0.003) 0.474 (- 0.006) 0.744 (- 0.014) 0.660 (- 0.016)
total average 0.601 (+0.000)
Table 19: General benchmark performance after supervised fine-tuning (SFT) on Thunder-NUBench with the detailed instruction. Here, acc denotes accuracy, acc_norm is normalized accuracy, and exact_match requires an exact string match with the reference answer. For each model and task, the main value reports the performance of the pre-SFT model, while the value in parentheses indicates the change after applying SFT.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate